# Efficient Computation of Voronoi Diagrams Using Point-in-Cell Tests

### v1

@@ 6.2. Nearest neighbor search
Figure 7. The numbers of point-in-cell tests invoked in Algorithm 2 using different versions of the edge-based bisector search.
[…]
By using different structures for nearest neighbor search, Algorithm 2 demonstrates varying levels of efficiency. The timing curves of Algorithm 2 using KD-tree and grids are shown in Fig. 8. The test domains include a rectangle, a surface sphere and a tetrahedralized ball, with sites randomly sampled at varying quantities. These curves indicate that Algorithm 2 achieves higher efficiency when using grids-based structure. Thus, we adopt the grids-based structure in both our serial and parallel algorithms.
6.3. Pre-clippings in parallel algorithm
In the implementation of our parallel algorithm, the structure for nearest neighbor searches resides in the GPU global memory, which is characterized by slow access speeds. Although the improved edge-based bisector search is applied, the number of nearest neighbor searches remains high in our method. Consequently, the algorithm requires frequent access to the GPU global memory, resulting in a significant time overhead. To further reduce the number of nearest neighbor searches in the parallel algorithm, we introduce pre-clippings prior to invoking Algorithm 3.
When computing the intersection $\mathcal{P}$ of a triangle/tetrahedron $\mathcal{T}$ and a Voronoi cell $\Omega_{i}$, an intermediate polygon/polytope $\mathcal{T}_{\text{temp}}$ that is smaller than $\mathcal{T}$ is first generated via pre-clippings, and Algorithm 3 then takes $\mathcal{T}_{\text{temp}}$ instead of $\mathcal{T}$ as input. The bisectors used in pre-clippings are those of site $\mathbf{v}_{i}$ and its several nearest neighbors, since they have a high probability to affect the result. Table 1 shows the improvement of our parallel algorithm with pre-clippings. After sufficient experiments, the optimal neighbor numbers of each site for pre-clippings under different cases are provided in Table 2, which are much less than those required in the kNN-based method [Ray et al., 2018; Liu et al., 2022].
@@ 6.3. Pre-clippings in parallel algorithm
Figure 8. The timing curves of Algorithm 2 using different structures for nearest neighbor search.
Table 1. Timings in seconds for computing 3D clipped Voronoi cells using our parallel algorithm with and without pre-clippings (cube, $m=1$).
configuration without pre-clippings with pre-clippings
$n$ 100K 500K 1M 100K 500K 1M
random sites 0.054 0.177 0.323 0.037 0.089 0.155
uniform sites 0.044 0.132 0.236 0.025 0.038 0.052
Table 2. Optimal neighbor numbers per site for pre-clippings.
case 2D surface 3D
random sites 12 25 35
grid sites 12 n/a 35
uniform sites 10 10 17
@@
DT-based method 0.003 0.113 0.588 0.003 0.103 0.517 0.003 0.11 0.532
kNN-based method 0.004 0.065 0.735 0.003 0.05 0.498 0.006 0.267 1.477
our method (Algorithm 2) 0.003 0.05 0.567 0.002 0.044 0.464 0.003 0.054 0.495
Figure 9. Clipped/restricted Voronoi diagrams of the given sites with different distributions. Top row: random sites; middle row: uniform sites; bottom row: uniform sites that are partially removed; left column: 2D, m=27,n=1m=27,n=1K; middle column: surface, m=20.8m=20.8K, n=5n=5K; right column: 3D, m=20.5m=20.5K, n=10n=10K. The time costs (in seconds) for each result of the DT-based, kNN-based methods, and our method (Algorithm 2) are provided in the inset table.
7. Results and Applications
@@
  • Random sites (white noise): The sites are randomly sampled from the given domain.
  • Uniform sites (blue noise): The sites are typically generated by centroidal Voronoi tessellation methods [Du et al., 1999; Liu et al., 2009].
  • Grid sites: The sites are sampled with random perturbations from regular grid patterns.
7.1. Serial performance
We compare our method with the DT-based and kNN-based methods while ensuring exact results. All these methods are implemented in C++, with the Delaunay triangulation for the DT-based method constructed using the CGAL library [The CGAL Project, 2022]. For the kNN-based method, the nearest neighbors list of each site is enlarged independently if the security radius is not met.
Fig. 9 shows a group of clipped or restricted Voronoi diagram results on 2D, surface, and 3D domains, with different site distributions, i.e., random sites (top row), uniform sites (middle row), and uniform sites that are partially removed (bottom row). The time costs of three methods for generating each result are given in the inset table. It is clear that the efficiency of the kNN-based method fluctuates more intensely than the DT-based method and ours. When a site dominates points far away from it, the kNN-based method has to launch many more clippings to meet the security radius for generating its cell, even though most of the clippings are invalid. Thus, it performs even worse than the DT-based method in such cases, see Fig. 9(c, g, h, i). In contrast, our method remains more stable for different site distributions, consumes much less time than the others, regardless of site distributions. Another group of results are given in Fig. 10, where our method still costs the least.
(a) 2D, m=180,n=1m=180,n=1K
(b) Surface, m=60.2m=60.2K, n=5n=5K
(c) 3D, m=15.6m=15.6K, n=10n=10K
Figure 10. Another group of results, where the site distribution is uniform. The time costs of the DT-based, kNN-based methods, and our method (Algorithm 2) are (a) 0.005, 0.005, 0.004; (b) 0.174, 0.09, 0.08; (c) 0.537, 0.51, 0.477 seconds, respectively.
@@ 7.1. Serial performance
Figure 11. Timing curves of the DT-based, kNN-based methods, and our method for computing clipped or restricted Voronoi diagrams across different domains with varying numbers of sites. The curves compare performance under random (top) and uniform (bottom) site distributions, respectively.
The timing curves of three methods against the number of sites are plotted in Fig. 11. The tested domains include: a rectangle (2D, $m=1$), two triangular surface meshes (sphere, m=1.2m=1.2K and bunny, m=60.2m=60.2K) and two tetrahedral meshes (ball, m=1.9m=1.9K and fertility, m=15.6m=15.6K). The site distributions correspond to random (top) and uniform (bottom) configurations, respectively. With increasing numbers of sites, the computational cost of Delaunay triangulation construction in the DT-based method grows significantly, resulting in the lowest efficiency in most cases. The kNN-based method performs the worst in 3D cases with random sites, while our method achieves the highest efficiency across all cases.
Furthermore, by fixing the number of sites (n=50n=50K), we compare the performance of three methods against the number of domain triangles/tetrahedra, their timing curves are presented in Fig. 12, where the domains are a triangular surface mesh (left) and a tetrahedral mesh (right), with site distributions being random (top) and uniform (bottom), respectively. This comparison reaffirms the superior performance of our method, since these curves exhibit a similar trend to that in Fig. 11: the kNN-based method is the most time-consuming in the 3D case with random sites, while the DT-based method performs worst in other scenarios. In contrast, our method consistently requires the least time.
@@
Figure 12. Timing curves of the DT-based, kNN-based methods, and our method for computing clipped or restricted Voronoi diagrams with varying numbers of domain triangles/tetrahedra, where the number of sites is fixed at n=50n=50K. The site distributions are random (top) and uniform (bottom).
7.2. Parallel performance
The parallel version of our method still guarantees exact results. Although the parallelized kNN-based method lacks such a guarantee, we continue to compare our method with it to demonstrate our parallel performance advantage. There are different parallel implementations of the kNN-based method in various domains: surface meshes [Han et al., 2017], regular volumes [Ray et al., 2018], irregular volumes [Liu et al., 2022]. The numbers of nearest neighbors per site used in these implementations are provided in Table 3, where the values are either suggested by the authors (where reported) or optimally selected through our tests. Notably, the implementation by Liu et al. [Liu et al., 2022] requires an additional parameter $k_{\text{tet}}$, defined as the number of nearest sites per tetrahedron. In the following experiments, to generate clipped Voronoi cells with a total volume exceeding $89\%$ of the domain volume, the values $k_{\text{tet}}$ are set to $140,400,700$ when n=100n=100K, 500500K, 11M, respectively.
@@ 7.2. Parallel performance
7.3. Computation of power diagrams
@@
The power diagram (weighted Voronoi diagram) [Aurenhammer, 1987] is an extension of the Voronoi diagram, and computing power diagrams as fast as possible is important to downstream applications. Given a set of weighted sites $\left\{(\mathbf{v}_{i},w_{i})\right\}_{i=1}^{n}$, a power diagram can be traditionally computed by using the regular triangulation (RT) based method [Edelsbrunner and Shah, 1996]. It has been proven that any power diagram equals to a restricted Voronoi diagram in higher one dimensional space [Xiao et al., 2023], so that the Voronoi generation methods can be directly used after the given weighted sites are lifted as $\left\{\mathbf{p}_{i}=(\mathbf{v}_{i},\sqrt{\eta-w_{i}})\right\}_{i=1}^{n}$, where $\eta\geq w_{\text{max}}=\max\left\{w_{1},\dots,w_{n}\right\}$. The kNN-based method has been extended to compute power diagrams [Xiao et al., 2023], and we apply our method to the computation in this section.
Fig. 13 shows the clipped or restricted power diagrams generated by our method, where the given weighted sites are with uniform locations and random weights, the weights are in $|\mathcal{M}|/n,|\mathcal{M}|/n]$ and $|\mathcal{M}|$ is the total area/volume of the given domain. To compare our method with the RT-based and kNN-based methods, we statistic the time costs of these methods to generate the same results. A similar conclusion is conducted that our method consumes the least computational time for each case. It is worth pointing out that because the weighted sites are lifted to a higher dimensional space, our method may consume more time on the nearest neighbor searches when computing power diagrams compared to computing Voronoi diagrams.
Table 3. The numbers of nearest neighbors per site used in parallel implementations of the kNN-based method.
case surface regular 2D/3D irregular 3D
random sites 70 70/190 90
grid sites n/a 25/90 n/a
uniform sites 12 10/35 35
It has been shown that the parameter $\eta$ influences the efficiency of the kNN-based method because it determines whether the security radius can be met early. Thus, the minimal $\eta$ (i.e., $\eta=w_{\text{max}}$) is optimal for the kNN-based method [Xiao et al., 2023]. Similarly, $\eta$ also has a high impact on our method’s efficiency. A larger $\eta$ generates lifted sites far away from the domain, causing our method to traverse more grids during each nearest neighbor search and thereby increasing computational time. Hence, we suggest using the minimum $\eta$ (i.e., $\eta=w_{\text{max}}$) for lifting weighted sites when computing power diagrams with our method.
Table 4. Timings in seconds for computing 2D clipped Voronoi cells (rectangle, $m=1$).
method DT-based (CPU) [Ray et al., 2018] (GPU) ours (Algorithm in 5.2, GPU)
$n$ 100K 500K 1M 100K 500K 1M 100K 500K 1M
random sites 0.758 18.979 70.921 0.019 0.039 0.063 0.015 0.023 0.03
grid sites 0.739 17.332 70.22 0.015 0.02 0.028 0.015 0.02 0.025
uniform sites 0.72 16.851 68.033 0.015 0.017 0.02 0.015 0.017 0.02
Table 5. Timings in seconds for computing restricted Voronoi cells (bunny, m=60.2m=60.2K).
method DT-based (CPU) [Han et al., 2017] (GPU) ours (Algorithm 4, GPU)
$n$ 100K 500K 1M 100K 500K 1M 100K 500K 1M
random sites 2.385 19.745 50.946 0.058 0.185 0.333 0.052 0.168 0.304
uniform sites 2.29 18.093 50.661 0.044 0.147 0.264 0.044 0.142 0.255
Table 6. Timings in seconds for computing 3D clipped Voronoi cells (cube, $m=1$).
method DT-based (CPU) [Ray et al., 2018] (GPU) ours (Algorithm in 5.2, GPU)
$n$ 100K 500K 1M 100K 500K 1M 100K 500K 1M
random sites 3.446 21.57 53.268 0.094 0.345 0.656 0.037 0.089 0.155
grid sites 2.553 20.456 52.886 0.049 0.147 0.265 0.035 0.072 0.118
uniform sites 2.439 20.355 51.377 0.028 0.046 0.063 0.025 0.038 0.052
Table 7. Timings in seconds for computing 3D clipped Voronoi cells (fertility, m=15.6m=15.6K).
method DT-based (CPU) [Liu et al., 2022] (GPU) ours (Algorithm 5, GPU)
$n$ 100K 500K 1M 100K 500K 1M 100K 500K 1M
random sites 4.374 27.395 66.287 0.122 0.404 0.755 0.096 0.256 0.442
uniform sites 4.056 26.845 65.143 0.066 0.17 0.3 0.048 0.107 0.175
7.4. Centroidal Voronoi tessellations
The centroidal Voronoi tessellation (CVT) is a special Voronoi diagram in which each site coincides with the centroid of its Voronoi cell. It also can be defined as the tessellation when the following energy function is minimized [Du et al., 1999; Liu et al., 2009]:
(4) $\mathcal{E}(\left\{\mathbf{v}_{i}\right\}_{i=1}^{n})=\sum_{i=1}^{n}\int_{\Omega_{i}}\rho(\mathbf{x})\|\mathbf{x}-\mathbf{v}_{i}\|^{2}d\mathbf{x},$
where $\rho(\mathbf{x})$ is the density function. Here, the Lloyd’s method [Du et al., 1999] is adopted to minimize equation (4). In each iteration, we compute centroids of Voronoi cells as the new sites and update the Voronoi diagram. Fig. 14 shows 2D and 3D CVT results after 200 iterations starting from random initialization, where $\rho(\mathbf{x})=e^{-20\|\mathbf{x}\|^{2}}+0.05\prod_{k=1}^{d}\sin^{2}(\pi\mathbf{x}(k))$, n=2n=2K and 10K in the 2D and 3D cases, respectively. The total time costs for computing centroids are 1.21 (2D) and 370.15 (3D) seconds. For the computation of Voronoi diagrams, we tested the DT-based, kNN-based methods and ours, which cost in total 0.779, 0.966, and 0.514 seconds in the 2D case, and 39.139, 85.298, and 28.064 seconds in the 3D case. Our method saves 34.02% (2D) / 28.3% (3D) of time compared to the DT-based method, and 46.79% (2D) / 67.1% (3D) of that compared to the kNN-based method, revealing the efficiency advantage of our method.
We also perform the Lloyd’s method on GPUs. Starting from Fig. 1(left) with 1M sites, the CVT result after 100 iterations is shown in Fig. 1(right), where $\rho(\mathbf{x})=1$. Our parallel algorithm and the parallelized kNN-based implementation [Ray et al., 2018] are both tested for generating Voronoi cells in this example. The result obtained using our parallel algorithm takes only 4.297 seconds, while it requires 6.215 seconds if the parallelized kNN-based implementation is applied.
8. Conclusion and Future Work
In this paper, we propose an efficient method for computing clipped or restricted Voronoi diagrams. The key distinction between our method and existing approaches lies in the generation of each triangle/tetrahedron-cell intersection. A clipping will be launched only if an edge of the intermediate intersection has endpoints being respectively inside and outside a target Voronoi cell. The corresponding clipping plane is rapidly identified through our presented edge-based bisector search scheme. Consequently, all the clippings in our method are valid to final results, demonstrating a great advantage over existing approaches. Additionally, we introduce a parallel implementation of our method, further accelerating Voronoi diagram computations. Extensive experimental evaluations confirm that our method outperforms state-of-the-art ones in both serial and parallel efficiency. It is worth exploring broader applications of our method in future work.
(a) 2D, m=1,n=1m=1,n=1K
(b) Surface, m=149.5m=149.5K, n=5n=5K
(c) 3D, m=24.5m=24.5K, n=10n=10K
@@ 8. Conclusion and Future Work
Figure 13. Our method can be directly applied to the computation of power diagrams after the given weighted sites are lifted. The given weighted sites are with uniform locations and random weights. The sites with positive and negative weights are rendered as red and blue disks/balls, respectively, and the absolute value of each weight is proportional to the size of corresponding disk/ball. To obtain these results, the time costs of the RT-based, kNN-based methods and ours are (a) 0.003, 0.002, 0.002, (b) 0.326, 0.166, 0.152, (c) 0.581, 0.572, 0.546 seconds, respectively.
(a) 2D, initialization
(b) 2D, result
(c) 3D, initialization
(d) 3D, result

### v2

@@ VI-B Nearest neighbor search
Figure 8: The numbers of point-in-cell tests invoked in Algorithm 2 using different versions of the edge-based bisector search.
Because each point-in-cell test only requires the nearest site of query point, a grids-based structure [36] is well-suited for this task. The given sites are first placed into a group of grids according to their coordinates. To find the nearest site for a query point, we examine nearby grids around the query point and traverse the sites within these grids to determine the closest one. More details can be found in [36]. We set the grid resolution so that each grid cell contains, on average, 2 $\sim$ 3 sites (for the 2D and surface cases) and 3 $\sim$ 4 sites (for the 3D case), respectively.
By using different structures for nearest neighbor searches, Algorithm 2 demonstrates varying levels of efficiency. The runtime curves of Algorithm 2 using KD-tree and grids, respectively, are shown in Fig. 9. The test domains include a rectangle, a surface sphere and a tetrahedralized ball, with sites randomly sampled at varying quantities. These curves indicate that Algorithm 2 achieves higher efficiency when using grids-based structure. Thus, we adopt the grids-based structure in both our serial and parallel algorithms.
VI-C 3D cell clipping on GPUs
To accommodate the GPU architecture, we employ the 3D cell data structure proposed by Ray et al. [36] when implementing Algorithm 5. This structure comprises an array of planes and an array of corner points, where each corner point is defined as the intersection of three planes. It can also be interpreted as a triangular mesh, i.e., a plane corresponds to a mesh vertex, and a corner point corresponds to a mesh triangle. Therefore, when clipping the simplex-cell intersection $\mathcal{P}_{a,i}$ of domain simplex $\mathcal{T}_{a}$ and Voronoi cell $\Omega_{i}$ using a bisector $B_{i,j}$, the procedure consists of the following steps: (1) identify all mesh triangles $R$ of $\mathcal{P}_{a,i}$ that lie outside $\Omega_{i}$; (2) determine the boundary loop of $R$; (3) remove $R$ and generate new mesh triangles by connecting $B_{i,j}$ to the boundary edges of $R$. For further details, refer to [36].
However, due to the finite precision of floating-point arithmetic, this cell-clipping procedure may misclassify the side of some mesh triangles and consequently generate a disconnected boundary for the set $R$. This issue cannot be avoided in the existing kNN-based parallel algorithms [36, 32, 4]. In contrast, Algorithm 3 readily avoids the problem. Because we always select a corner point $\mathbf{p}$ that is guaranteed to lie outside $\Omega_{i}$ for each clipping (see Fig. 7(a, c)), we simply replace $R$ with $R_{\mathbf{p}}$ in the steps above, where $R_{\mathbf{p}}$ denotes the connected component of $R$ that contains $\mathbf{p}$.
@@
Figure 9: The runtime curves of Algorithm 2 using different structures for nearest neighbor searches.
@@
DT-based method [2, 46, 47] 0.003 0.113 0.6 0.003 0.103 0.48 0.003 0.11 0.49
kNN-based method [30] 0.004 0.065 0.684 0.003 0.05 0.465 0.006 0.267 1.296
our method (Algorithm 2) 0.003 0.05 0.462 0.002 0.044 0.403 0.003 0.054 0.433
Figure 10: Clipped/restricted Voronoi diagrams of the given sites with different distributions. Top row: white noise; middle row: blue noise; bottom row: blue noise with partial sites removal; left column: 2D, m=27,n=1m=27,n=1K; middle column: surface, m=20.8m=20.8K, n=5n=5K; right column: 3D, m=20.5m=20.5K, n=10n=10K. The runtimes (in seconds) for each result of the DT-based [2, 46, 47], kNN-based [30] methods, and our method (Algorithm 2) are provided in the inset table.
(a) m=180,n=1m=180,n=1K
(b) m=60.2m=60.2K, n=5n=5K
(c) m=15.6m=15.6K, n=10n=10K
Figure 11: Another group of results, where the sites are with blue noise. The runtimes of the DT-based [2, 46, 47], kNN-based [30] methods, and our method (Algorithm 2) are (a) 2D: 0.005, 0.005, 0.004; (b) surface: 0.125, 0.09, 0.08; (c) 3D: 0.537, 0.51, 0.39 seconds, respectively.
VII Results and Applications
@@
  • White noise: The sites are randomly sampled from the given domain.
  • Blue noise: The sites are typically generated by centroidal Voronoi tessellation methods [17, 33].
  • Perturbed grid: The sites are sampled with random perturbations from regular grid patterns.
@@
Figure 12: Runtime curves of the DT-based [2, 46, 47], kNN-based [30] methods, and our method (Algorithm 2) for computing clipped or restricted Voronoi diagrams across different domains with varying numbers of sites. The curves compare performance under white noise (top) and blue noise (bottom), respectively.
VII-A Serial performance
We compare our method with the DT-based [2, 46, 47] and kNN-based [30] methods while ensuring exact results. All these methods are implemented in C++, with the Delaunay triangulation for the DT-based method constructed using the CGAL library [41]. For the kNN-based method, the nearest neighbors list of each site is enlarged independently if the security radius is not met.
Fig. 10 presents clipped or restricted Voronoi diagrams on 2D, surface, and 3D domains. The rows correspond to different site distributions: white noise (top), blue noise (middle), and blue noise with partial sites removal (bottom). The runtimes of three methods for generating each result are given in the inset table. It is clear that the efficiency of the kNN-based method fluctuates more intensely than the DT-based method and ours. When a site dominates points far away from it, the kNN-based method has to launch many more clippings to meet the security radius for generating its cell, even though most of the clippings are invalid. Thus, it performs even worse than the DT-based method in such cases, see Fig. 10(c, g, h, i). In contrast, our method remains more stable and consumes much less time than the others across all site distributions. Another group of results are given in Fig. 11, where our method still costs the least.
The runtime curves of three methods against the number of sites are plotted in Fig. 12. The tested domains include: a rectangle (2D, $m=1$), two triangular surface meshes (sphere, m=1.2m=1.2K and bunny, m=60.2m=60.2K) and two tetrahedral meshes (ball, m=1.9m=1.9K and fertility, m=15.6m=15.6K). The site distributions correspond to white noise (top) and blue noise (bottom), respectively. With increasing numbers of sites, the computational cost of Delaunay triangulation construction in the DT-based method grows significantly, resulting in the lowest efficiency in most cases. The kNN-based method performs the worst in the 3D case with white noise, while our method achieves the highest efficiency across all cases.
@@ VII-A Serial performance
Figure 13: Runtime curves of the DT-based, kNN-based methods, and our method for computing clipped or restricted Voronoi diagrams with varying numbers of domain simplices, where the number of sites is fixed at n=50n=50K. The site distributions are white noise (top) and blue noise (bottom).
Furthermore, by fixing the number of sites (n=50n=50K), we compare the performance of three methods against the number of domain simplices, their runtime curves are presented in Fig. 13, where the domains are a triangular mesh (left) and a tetrahedral mesh (right), and the site distributions are white noise (top) and blue noise (bottom), respectively. This comparison reaffirms the superior performance of our method, since these curves exhibit a similar trend to that in Fig. 12: the kNN-based method is the most time-consuming in the 3D case with randomly generated sites, while the DT-based method performs worst in other scenarios. In contrast, our method consistently consumes the least time.
(a) white noise
(b) blue noise
Figure 14: Runtime comparison between Han et al. [26] and Algorithm 4 on a surface domain (bunny, m=60.2m=60.2K) with varying site numbers. The sites are with white noise (a) and blue noise (b), respectively.
(a) white noise
(b) perturbed grid
(c) blue noise
Figure 15: Runtime comparison between Ray et al. [36], Basselin et al. [4] and our parallel algorithm (Section V-B) on a simple 3D domain (cube, $m=1$) with varying site numbers. The sites are with white noise (a), perturbed grid (b) and blue noise (c), respectively.
(a) white noise
(b) blue noise
Figure 16: Runtime comparison between Liu et al. [32], Basselin et al. [4] and Algorithm 5 on a complex 3D domain (fertility, m=15.6m=15.6K) with varying site numbers. The sites are with white noise (a) and blue noise (b), respectively.
(a) white noise
(b) perturbed grid
(c) blue noise
Figure 17: Comparison of peak GPU memory usage between Ray et al. [36], Basselin et al. [4] and our parallel algorithm (Section V-B) on a simple 3D domain (cube, $m=1$) with varying site numbers. The sites are with white noise (a), perturbed grid (b) and blue noise (c), respectively.
VII-B Parallel performance
The parallel version of our method still ensures exact results and rapidly delivers outcomes under varying site distributions. To evaluate the performance of our parallel algorithm, we compare it with the state-of-the-art method, namely the kNN-based method. Several parallel versions of the kNN-based method have been developed for different domains: surface meshes [26], simple 3D volumes [36], and complex 3D volumes [32, 4]. We implement our parallel algorithm and the algorithm of Han et al. [26] using CUDA, while the implementations of other kNN-based versions are obtained from the following sources: Ray et al. [36] [^1] https://dl.acm.org/doi/10.1145/3272127.3275092, Liu et al. [32] [^2] https://github.com/xh-liu-tech/3D-Voronoi-GPU, and Basselin et al.  [4] [^3] https://github.com/basselin7u/GPU-Restricted-Power-Diagrams. All these algorithms are tested using 32-bit floating-point numbers.
@@ VII-B Parallel performance
Our parallel algorithm not only improves efficiency but also reduces GPU memory requirements. The parallel implementations of the kNN-based method must store the k-nearest neighbors for each site, resulting in substantial memory overhead. Although Basselin et al.  [4] introduced a multi-pass strategy to alleviate this demand, memory consumption remains high. Our algorithm, however, only requires a structure for nearest-neighbor searches, leading to markedly reduced GPU memory requirements. Fig. 17 compares the peak GPU memory usage of our parallel algorithm with those of Ray et al. [36] and Basselin et al.  [4] on a cube under varying site distributions and scales. The results show that our algorithm consumes significantly less memory across all tested cases.
VII-C Computation of power diagrams
@@
The power diagram (weighted Voronoi diagram) [1] is an extension of the Voronoi diagram, and computing power diagrams as fast as possible is important to downstream applications. Given a set of weighted sites $\left\{(\mathbf{v}_{i},w_{i})\right\}_{i=1}^{n}$, a power diagram can be traditionally computed by using the regular triangulation (RT) based method [20]. It has been proven that any power diagram equals to a restricted Voronoi diagram in higher one dimensional space [43], so that the Voronoi generation methods can be directly used after the given weighted sites are lifted as $\left\{\mathbf{p}_{i}=(\mathbf{v}_{i},\sqrt{\eta-w_{i}})\right\}_{i=1}^{n}$, where $\eta\geq w_{\text{max}}=\max\left\{w_{1},\dots,w_{n}\right\}$. The kNN-based method has been extended to compute power diagrams [43], and we apply our method to the computation in this section.
(a) m=1,n=1m=1,n=1K
(b) m=149.5m=149.5K, n=5n=5K
(c) m=24.5m=24.5K, n=10n=10K
@@
Figure 18: Our method can be directly applied to the computation of power diagrams after the given weighted sites are lifted. The given weighted sites are with blue noise locations and random weights. The sites with positive and negative weights are rendered as red and blue disks/balls, respectively, and the absolute value of each weight is proportional to the size of corresponding disk/ball. To obtain these results, the runtimes of the RT-based, kNN-based methods and ours are (a) 2D: 0.003, 0.002, 0.002, (b) surface: 0.326, 0.166, 0.145, (c) 3D: 0.552, 0.526, 0.46 seconds, respectively.
Fig. 18 shows three examples of clipped or restricted power diagrams generated by our method, where the given weighted sites are with blue noise locations and random weights, the weights are in $|\mathcal{M}|/n,|\mathcal{M}|/n]$ and $|\mathcal{M}|$ is the total area/volume of the given domain. We use the RT-based and kNN-based methods and our method to generate the same results and compare their running time. A similar conclusion is conducted that our method consumes the least computational time for each case. It is worth pointing out that because the weighted sites are lifted to a higher dimensional space, our method may consume more time on the nearest neighbor searches when computing power diagrams compared to computing Voronoi diagrams.
It has been shown that the parameter $\eta$ influences the efficiency of the kNN-based method because it determines whether the security radius can be met early. Thus, the minimal $\eta$ (i.e., $\eta=w_{\text{max}}$) is optimal for the kNN-based method [43]. Similarly, $\eta$ also has a high impact on our method’s efficiency. A larger $\eta$ generates lifted sites far away from the domain, causing our method to traverse more grids during each nearest neighbor search and thereby increasing computational time. Hence, we suggest using the minimal $\eta$ (i.e., $\eta=w_{\text{max}}$) for lifting weighted sites when computing power diagrams with our method.
VII-D Centroidal Voronoi tessellations
The centroidal Voronoi tessellation (CVT) is a special Voronoi diagram in which each site coincides with the centroid of its Voronoi cell. It also can be defined as the tessellation when the following energy function is minimized [17, 33]:
$\mathcal{E}(\left\{\mathbf{v}_{i}\right\}_{i=1}^{n})=\sum_{i=1}^{n}\int_{\Omega_{i}}\rho(\mathbf{x})\|\mathbf{x}-\mathbf{v}_{i}\|^{2}d\mathbf{x},$ (3)
where $\rho(\mathbf{x})$ is the density function. Here, the Lloyd’s method [17] is adopted to minimize equation (3). In each iteration, we compute centroids of Voronoi cells as the new sites and update the Voronoi diagram. Fig. 19 shows 2D and 3D CVT results after 200 iterations starting from random initialization, where $\rho(\mathbf{x})=e^{-20\|\mathbf{x}\|^{2}}+0.05\prod_{k=1}^{d}\sin^{2}(\pi\mathbf{x}(k))$, n=2n=2K and 10K in the 2D and 3D cases, respectively. The total time costs for computing centroids are 1.21 (2D) and 370.15 (3D) seconds. For the computation of Voronoi diagrams, we tested the DT-based [2, 47], kNN-based [30] methods and ours, which cost in total 0.779, 0.966, and 0.514 seconds in the 2D case, and 39.139, 85.298, and 22.535 seconds in the 3D case. Our method saves 34% (2D) / 42.4% (3D) of time compared to the DT-based method, and 46.8% (2D) / 73.5% (3D) of that compared to the kNN-based method, revealing the efficiency advantage of our method.
We also perform the Lloyd’s method on GPUs. Starting from Fig. 1(a) with 5M sites distributed in white noise, the CVT result after 100 iterations is shown in Fig. 1(b), where $\rho(\mathbf{x})=1$. Our parallel algorithm and the kNN-based parallel algorithm [36] are both tested for generating Voronoi cells in this example. The result obtained using our parallel algorithm takes only 31.07 seconds, while it requires 54.8 seconds if the kNN-based parallel algorithm [36] is applied.
2D CVT 3D CVT
Refer to caption Refer to caption Refer to caption Refer to caption
initialization result initialization result